# Fan distributions via Tverberg partitions and Gale duality

### v2

@@ 3.1. Real Linear Fans
Gale duality gives an equivalence between proper $r$-Tverberg partitions of sequences of points in $\mathbb{R}^{d}$ and distributions of sequences of points by linear $r$-fans of codimension $r-2$.
@@
    Suppose that $(\sigma_{1},\ldots,\sigma_{r})$ is a proper $r$-Tverberg tuple for $f$ and let $I_{1},\ldots,I_{r}\subset[n]$ be the corresponding vertex sets of the $\sigma_{j}$. Then there exists a linear $r$-fan $F_{r}$ in $\mathbb{R}^{n-d-1}$ of codimension $r-2$ which distributes the $g_{i}$ according to the $I_{j}$.
@@
    Suppose that $I_{1},\ldots,I_{r}$ are non-empty pairwise disjoint subsets of $[n]$ and that $F_{r}$ is an $r$-fan in $\mathbb{R}^{n-d-1}$ of codimension $r-2$ which distributes the $g_{i}$ according to the $I_{j}$. Then there exists a proper $r$-Tverberg tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f$ such that $\operatorname{\mathrm{Vert}}(\sigma_{j})=I_{j}$ for all $j\in[r]$.
@@
$\langle\alpha_{j},g_{i}\rangle=(\lambda_{j})_{i}$ (3.1)
@@
We claim that any $r-1$ of the $\alpha_{j}$ are linearly independent, but that $\alpha_{1},\ldots,\alpha_{r}$ are not. To see this, observe that since the $g_{i}$ linearly span $\mathbb{R}^{n-d-1}$, equation (3.1) implies that any linear relation $\sum_{j=1}^{n-d-1}\mu_{j}\alpha_{j}=0$ holds among the $\alpha_{j}$ if and only if the corresponding relation $\sum_{j=1}^{n-d-1}\mu_{j}\lambda_{j}=0$ holds among the $\lambda_{j}$. It is easily seen that any $r-1$ vectors among the $\lambda_{j}$ are linearly independent, so any $r-1$ of the $\alpha_{j}$ are as well. Similarly, $\sum_{j=1}^{r}\lambda_{j}=0$ implies that $\sum_{j=1}^{r}\alpha_{j}=0$. For each $j\in[r]$, let $H_{j}=\langle\alpha_{j}\rangle^{\perp}$ be the linear hyperplane corresponding to $\alpha_{j}$. Now let $A_{j}=\cap_{i\neq j,j-1}H_{i}$ and $B_{j}=H_{j}^{+}\cap A_{j}$, where $H_{j}^{+}=\{x\in\mathbb{R}^{n-d-1}\mid\langle\alpha_{j},x\rangle\geq 0\}$. Thus $H_{j}\cap A_{j}=\cap_{i=1}^{r}H_{i}$ for all $j\in[r]$ and $F_{r}=\cup_{j=1}^{r}B_{j}$ is a linear $r$-fan of codimension $r-2$. The distribution claim about the $g_{i}$ also follows from equation (3.1). For if $j\in[r]$ and $i\in I_{j}$, we have $\langle\alpha_{j},g_{i}\rangle=(\lambda_{j})_{i}=t_{i}>0$, while $\langle\alpha_{k},g_{i}\rangle=(\lambda_{k})_{i}=0$ if $k\neq j-1,j$. Thus $g_{i}\in\operatorname{\mathrm{Int}}(B_{j})=\operatorname{\mathrm{Int}}(H_{j}^{+})\cap A_{j}$. On the other hand, $\langle g_{i},\alpha_{j}\rangle=0$ for all $j\in[r]$ if $i\notin\sqcup_{j=1}^{r}I_{j}$ and therefore $g_{i}\in C$.
@@
For part (b), suppose that $F_{r}$ a linear $r$-fan in $\mathbb{R}^{n-d-1}$ of codimension $r-2$ determined by $H_{1}=\langle\alpha_{1}\rangle^{\perp},\ldots,H_{r}=\langle\alpha_{r}\rangle^{\perp}$ in $\mathbb{R}^{n-d-1}$. By assumption, $\alpha_{1},\ldots,\alpha_{r}$ are linearly dependent, so $\sum_{j=1}^{r}\mu_{j}\alpha_{j}=0$ for some scalars $\mu_{1},\ldots,\mu_{r}$, not all of which are zero. Since any $r-1$ of the $\alpha_{j}$ are linearly independent, $\mu_{j}\neq 0$ for all $j\in[r]$. Thus by scaling both $\alpha_{j}$ and $\beta_{j}$ by $\mu_{j}$ we may assume that $\sum_{j=1}^{r}\alpha_{j}=0$. Without loss of generality, we may let $H^{+}_{j}=\{x\in\mathbb{R}^{n-d-1}\mid\langle\alpha_{j},x\rangle\geq 0\}$ for all $j\in[r]$.
@@
For each $j\in[r]$, define $\lambda_{j}=(\lambda_{j,1},\ldots,\lambda_{j,n})\in\mathbb{R}^{n}$ by $\lambda_{j,i}=\langle\alpha_{j},g_{i}\rangle$ for all $i\in[n]$. As $g_{i}\in\operatorname{\mathrm{Int}}(B_{j})$ for all $i\in I_{j}$, $\lambda_{j}$ is non-zero and so by Proposition 2.2 we have $\sum_{i=1}^{n}\lambda_{j,i}=0$ and $\sum_{i=1}^{n}\lambda_{j,i}a_{i}=0$. Now let $j\in[r]$ and suppose that $i\in I_{j}$. Since $g_{i}\in\operatorname{\mathrm{Int}}(B_{j})$, we have $\lambda_{j,i}>0$ and $\lambda_{k,i}=0$ for all $k\neq j-1,j$. As $\sum_{j=1}^{r}\alpha_{j}=0$, we have $\langle\sum_{j=1}^{r}\alpha_{j},g_{i}\rangle=\langle\alpha_{j-1},g_{i}\rangle+\langle\alpha_{j},g_{i}\rangle=0$. Thus $\lambda_{j-1,i}=-\lambda_{j,i}$ for all $i\in I_{j}$. Again fix $j\in[r]$ and consider the relation $\sum_{i=1}^{n}\lambda_{j,i}a_{i}=0$. If $i\notin I_{j}\sqcup I_{j+1}$, then $g_{i}\notin\operatorname{\mathrm{Int}}(B_{j})\sqcup\operatorname{\mathrm{Int}}(B_{j+1})$. In particular, $g_{i}\in\cap_{\ell\neq k,k-1}H_{\ell}$ for some $k\neq j,j+1$ and so $\langle\alpha_{j},g_{i}\rangle=0$ for all $i\notin I_{j}\sqcup I_{j+1}$. Thus $\sum_{i\in I_{j}}\lambda_{j,i}a_{i}+\sum_{i\in I_{j+1}}\lambda_{j,i}a_{i}=0$ and so $\sum_{i\in I_{j}}\lambda_{j,i}a_{i}=-\sum_{i\in I_{j+1}}\lambda_{j,i}a_{i}=\sum_{i\in I_{j+1}}\lambda_{j+1,i}a_{i}$ for any $j\in[r]$. We therefore have
$\sum_{i\in I_{1}}\lambda_{i}a_{i}=\cdots=\sum_{i\in I_{r}}\lambda_{i}a_{i}.$ (3.2)
@@ 3.2. Complex Linear Fans
Before giving a complex analogue of Lemma 3.1, we first give a description of complex regular $r$-fans in $\mathbb{C}^{d}$ in terms of the standard Hermitian inner product. Any complex hyperplane in $\mathbb{C}^{d}$ is of the form $H_{\mathbb{C}}:=\{z\in\mathbb{C}^{d}\mid\langle\alpha,z\rangle=\beta\}$ for some $(\alpha,\beta)\in\mathbb{C}^{n}\setminus\{0\}\times\mathbb{C}$. Let $\omega_{r}=\exp(2\pi i/r)$ be the standard $r$-th root of unity. If $F_{r}=\cup_{j=1}^{r}B_{j}$ is a complex regular $r$-fan centered about $H_{\mathbb{C}}$, then there exists some unit complex number $\lambda$ such that $B_{j}=\{z\in\mathbb{C}^{d}\mid\langle\alpha,z\rangle=\beta+t\lambda\omega_{r}^{j}\,\,\text{for some}\,\,t>0\}$ for all $j\in[r]$. Moreover, one may take $\lambda=1$ after scaling $(\alpha,\beta)$ by $\lambda^{-1}$.
@@
As before, suppose that $I_{1},\ldots,I_{r}\subset[n]$ are pairwise disjoint non-empty sets such that $\sum_{i\in I_{1}}t_{i}a_{i}=\cdots=\sum_{i\in I_{r}}t_{i}a_{i}$ where $t_{i}>0$ for all $i\in\sqcup_{j=1}^{r}I_{j}$ and $\sum_{i\in I_{j}}t_{i}=1$ for all $j\in[r]$. To construct a complex affine dependence among the $a_{i}$, define $\lambda=(\lambda_{1},\ldots,\lambda_{n})\in\mathbb{C}^{n}\setminus\{0\}$ as follows: for each $j\in[r]$ and each $i\in I_{j}$, let $\lambda_{i}=t_{i}\omega_{r}^{j}$, and let $\lambda_{i}=0$ for all $i\notin\sqcup_{j=1}I_{j}$. As $\sum_{i\in I_{j}}t_{j}=1$ for all $j\in[r]$, we have $\sum_{i=1}^{n}\lambda_{i}=\sum_{j\in[r]}\sum_{i\in I_{j}}t_{i}\omega_{r}^{j}=\sum_{j\in[r]}\omega_{j}^{r}=0$. Likewise, we have $\sum_{i\in I_{j}}t_{i}a_{i}=c$ is constant over $j\in[r]$ and so $\sum_{i\in[n]}\lambda_{i}a_{i}=\sum_{j\in[r]}\sum_{i\in I_{j}}t_{i}\omega_{r}^{j}a_{i}=c\sum_{j\in[r]}\omega_{r}^{j}=0$ as well. By Proposition 2.2, there is some $\alpha\neq 0$ such that $\langle\alpha,g_{i}\rangle=\lambda_{i}$ for all $i\in[n]$. Let $H_{\mathbb{C}}=\langle\alpha\rangle^{\perp}$ be the resulting linear complex hyperplane in $\mathbb{C}^{n-d-1}$, and for each $j\in[r]$ let $B_{j}=\{z\in\mathbb{C}^{n-d-1}\mid\langle\alpha,z\rangle=t\omega_{r}^{j}\,\,\text{for some}\,\,t\geq 0\}$. Thus $F_{r}=\cup_{j=1}^{r}B_{j}$ is a complex regular $r$-fan $F_{r}$ centered about $H_{\mathbb{C}}$. By construction, $g_{i}\in\operatorname{\mathrm{Int}}{B_{j}}$ for all $i\in I_{j}$ and any $j\in[r]$, while $g_{i}\in H_{\mathbb{C}}$ if $i\notin\sqcup_{j\in[r]}I_{j}$. ∎
@@ 4. Distributions by a Single Fan
We now prove our distribution results for a single fan. Each of these follows by applying Gale duality to the linear versions of associated topological Tverberg-type theorem. We defer the proof of the near optimality of Theorem of the real case of Theorem 1.4) Section 5.2. Analogous remarks apply to the complex case of Theorem 1.3 when $r=2$.
4.1. Proof of Theorems 1.11.3, and  1.4
@@
Let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. We let $n\geq(r-1)(d+m+1)$ when $\mathbb{K}=\mathbb{R}$ and $n\geq(r-1)(2d+m+1)+1$ when $\mathbb{K}=\mathbb{C}$. Suppose that $X=X_{1}\sqcup\cdots\sqcup X_{m}$ is an $m$-coloring of a set of $n$ affinely spanning points in $\mathbb{K}^{n-d-1}$. Corresponding to each $X_{i}$, define a corresponding family of non-empty subsets of $X$ by $\mathcal{F}_{i}=\{F\subset X\mid|F\cap X_{i}|>|X_{i}|/r\}$ and let $\mathcal{F}=\cup_{i=1}^{m}\mathcal{F}_{i}$. As no $\mathcal{F}_{i}$ can contain $r$ pairwise disjoint members by the pigeon-hole principle, $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$. Theorem 1.4(a) guarantees an $r$-fan $F_{r}=\cup_{j=1}^{r}B_{j}$ in $\mathbb{K}^{n-d-1}$ (an $r$-fan of codimension $r-2$ when $\mathbb{K}=\mathbb{R}$, and a complex regular fan when $\mathbb{K}=\mathbb{C})$ such that no $\operatorname{\mathrm{Int}}(B_{j})$ contains any element of $\mathcal{F}$. It follows that $|\operatorname{\mathrm{Int}}(B_{j})\cap X_{i}|\leq|X_{i}|/m$ for every $j\in[r]$ and each $i\in[m]$, so that $F_{r}$ is the desired $r$-fan. Indeed, were $|\operatorname{\mathrm{Int}}(B_{j})\cap X_{i}|>|X_{i}|/m$ for some $j\in[r]$ and $i\in[m]$ we would have $\operatorname{\mathrm{Int}}(B_{j})\cap X\in\mathcal{F}_{i}$, and therefore that $\operatorname{\mathrm{Int}}(B_{j})\cap X$ is not a subset of $\operatorname{\mathrm{Int}}(B_{j})$. ∎
We now prove Theorem 1.4.
@@
Again let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$ and let $X=\{x_{1},\ldots,x_{n}\}$ be a set of $n$ affinely spanning points in $\mathbb{K}^{n-d-1}$, where $n\geq(r-1)(d+m+1)+1$ when $\mathbb{K}=\mathbb{R}$ and $n\geq(r-1)(2d+m+1)+1$ if $\mathbb{K}=\mathbb{C}$. We are given that $\mathcal{F}$ is a family of non-empty subsets of $X$ satisfying $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$. We now lift $X$ to the $\mathbb{K}$-hyperplane $\mathbb{K}^{n-d-1}\times\{1\}$ of height one in $\mathbb{K}^{n-d}$. The sequence $g_{1}=(x_{1},1),\ldots,g_{n}=(x_{n},1)$ in $\mathbb{K}^{n-d}$ is linearly spanning, and including $g_{n+1}=-\sum_{i=1}^{n}g_{i}$ gives a sequence of $n+1$ points in $\mathbb{K}^{n-d}$ which both linearly span and which sum to zero. By Proposition 2.1, the sequence $g_{1},\ldots,g_{n+1}$ is the Gale transform of $n+1$ points $a_{1},\ldots,a_{n+1}$ in $\mathbb{K}^{d}$. We let $f\colon\Delta_{n}\rightarrow\mathbb{K}^{d}$ be the (real) linear map determined by sending each $i\in[n+1]$ to $a_{i}$ (here we identify $\mathbb{C}^{d}$ with $\mathbb{R}^{2d}$).
@@
Corresponding to $\mathcal{F}$, let $\mathcal{F}^{\prime}$ be the family of all subsets $\{i_{1},\ldots,i_{k}\}$ of $[n]$ such that $\{x_{i_{1}},\ldots,x_{i_{k}}\}$ is in $\mathcal{F}$. As $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$, $\chi(\mathrm{KG}^{r}(\mathcal{F}^{\prime}))\leq m$ as well. Given the value of $n$, we may apply Theorem 1.5 to the restriction $f\mid_{\Delta_{n-1}}\colon\Delta_{n-1}\rightarrow\mathbb{K}^{d}$ of $f$ to the simplex $\Delta_{n-1}$ whose vertex set is $[n]$ Thus there is a proper Tverberg pair $(\sigma_{1},\ldots,\sigma_{r})$ for $f|_{\Delta_{n-1}}$ such that each $I_{j}=\operatorname{\mathrm{Vert}}(\sigma_{j})\subset[n]$ does not contain any element of $\mathcal{F}^{\prime}$. As $(\sigma_{1},\ldots,\sigma_{r})$ is also a proper $r$-Tverberg partition for the original map $f$, by Lemma 3.1 (respectively, Lemma 3.2) there is a linear $r$-fan $F_{r}$ in $\mathbb{K}^{n-d}$ (of codimension $r-2$ when $r=2$ and complex linear when $\mathbb{K}=\mathbb{C}$) which distributes $g_{1},\ldots,g_{n+1}$ according to the $I_{j}$. Thus $g_{i}\in\operatorname{\mathrm{Int}}(B_{j})$ for all $i\in I_{j}$ and all $j\in[r]$, while $g_{i}\in C=\cap_{j=1}^{r}H_{j}$ for all $i\notin\sqcup_{j=1}^{r}I_{j}$. As $n+1$ does not lie in any of the $I_{j}$, we have that $g_{n+1}$ lies on $C$.
Now let $\mathbb{K}=\mathbb{R}$. For each $j\in[r]$ we have $H_{j}=\langle\alpha_{j}\rangle^{\perp}$ where $\alpha_{j}=(\beta_{j},\gamma_{j})\in\mathbb{R}^{n-d-1}\times\mathbb{R}$ is non-zero. The $r$-fan is of the form $F_{r}=\cup_{j=1}^{r}B_{j}$, where each $B_{j}$ is of the form $B_{j}=H_{j}^{+}\cap A_{j}$ with $A_{j}=\cap_{i\neq j,j-1}H_{j}$ and $H_{j}^{+}$ being some open half-space of $H_{j}^{+}$. For each $j\in[r]$, let $H^{\prime}_{j}=H_{j}\cap(\mathbb{R}^{n-d-1}\times\{1\})$ be the intersection of $H_{j}$ with the plane of height one in $\mathbb{R}^{n-d}$. Thus $H^{\prime}_{j}=\{(x,1)\in\mathbb{R}^{n-d-1}\times\{1\}\mid\langle\beta_{j},x\rangle=-\gamma_{j}\}$. Likewise, let $B^{\prime}_{j}=B_{j}\cap(\mathbb{R}^{n-d-1}\times\{1\})$ for all $j\in[r]$. Letting $H_{j}^{\prime\prime}$ and $B_{j}^{\prime\prime}$ be the corresponding projections of $H_{j}^{\prime}$ and $B_{j}^{\prime}$ onto $\mathbb{R}^{n-d-1}$, we will show that $F^{\prime\prime}_{r}=\cup_{j=1}^{r}B_{j}^{\prime\prime}$ is an $r$-fan of codimension $r-2$ in $\mathbb{R}^{n-d-1}$.

### v3

@@ 3.1. Linear Conical Fans
Gale duality gives an equivalence between proper $r$-Tverberg partitions of sequences of points in $\mathbb{R}^{d}$ and distributions of sequences of points by linear conical $r$-fans.
@@
    Suppose that $(\sigma_{1},\ldots,\sigma_{r})$ is a proper $r$-Tverberg tuple for $f$ and let $I_{1},\ldots,I_{r}\subset[n]$ be the corresponding vertex sets of the $\sigma_{j}$. Then there exists a linear conical $r$-fan $F_{r}$ in $\mathbb{R}^{n-d-1}$ which distributes the $g_{i}$ according to the $I_{j}$.
@@
    Suppose that $I_{1},\ldots,I_{r}$ are non-empty pairwise disjoint subsets of $[n]$ and that $F_{r}$ is a linear conical $r$-fan in $\mathbb{R}^{n-d-1}$ which distributes the $g_{i}$ according to the $I_{j}$. Then there exists a proper $r$-Tverberg tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f$ such that $\operatorname{\mathrm{Vert}}(\sigma_{j})=I_{j}$ for all $j\in[r]$.
@@
(3.1) $\langle\alpha_{j},g_{i}\rangle=(\lambda_{j})_{i}$
@@
We claim that any r1r-1 of the $\alpha_{j}$ are linearly independent, but that $\alpha_{1},\ldots,\alpha_{r}$ are not. To see this, observe that since the $g_{i}$ linearly span $\mathbb{R}^{n-d-1}$, equation (3.1) implies that any linear relation $\sum_{j=1}^{n-d-1}\mu_{j}\alpha_{j}=0$ holds among the $\alpha_{j}$ if and only if the corresponding relation $\sum_{j=1}^{n-d-1}\mu_{j}\lambda_{j}=0$ holds among the $\lambda_{j}$. It is easily seen that any r1r-1 vectors among the $\lambda_{j}$ are linearly independent, so any r1r-1 of the $\alpha_{j}$ are as well. Similarly, $\sum_{j=1}^{r}\lambda_{j}=0$ implies that $\sum_{j=1}^{r}\alpha_{j}=0$. For each $j\in[r]$, let $H_{j}=\langle\alpha_{j}\rangle^{\perp}$ be the linear hyperplane corresponding to $\alpha_{j}$. Now let $A_{j}=\cap_{i\neq j,j-1}H_{i}$ and $B_{j}=H_{j}^{+}\cap A_{j}$, where $H_{j}^{+}=\{x\in\mathbb{R}^{n-d-1}\mid\langle\alpha_{j},x\rangle\geq 0\}$. Thus $H_{j}\cap A_{j}=\cap_{i=1}^{r}H_{i}$ for all $j\in[r]$ and $F_{r}=\cup_{j=1}^{r}B_{j}$ is a linear conical $r$-fan. The distribution claim concerning the $g_{i}$ also follows from equation (3.1). For if $j\in[r]$ and $i\in I_{j}$, we have $\langle\alpha_{j},g_{i}\rangle=(\lambda_{j})_{i}=t_{i}>0$, while $\langle\alpha_{k},g_{i}\rangle=(\lambda_{k})_{i}=0$ if $k\neq j-1,j$. Thus $g_{i}\in\operatorname{\mathrm{Int}}(B_{j})=\operatorname{\mathrm{Int}}(H_{j}^{+})\cap A_{j}$. On the other hand, $\langle g_{i},\alpha_{j}\rangle=0$ for all $j\in[r]$ if $i\notin\sqcup_{j=1}^{r}I_{j}$ and therefore $g_{i}\in C$.
@@
For part (b), suppose that $F_{r}$ is a linear conical $r$-fan in $\mathbb{R}^{n-d-1}$ determined by the hyperplanes $H_{1}=\langle\alpha_{1}\rangle^{\perp},\ldots,H_{r}=\langle\alpha_{r}\rangle^{\perp}$ in $\mathbb{R}^{n-d-1}$. By assumption, $\alpha_{1},\ldots,\alpha_{r}$ are linearly dependent, so $\sum_{j=1}^{r}\mu_{j}\alpha_{j}=0$ for some scalars $\mu_{1},\ldots,\mu_{r}$, not all of which are zero. Since any r1r-1 of the $\alpha_{j}$ are linearly independent, $\mu_{j}\neq 0$ for all $j\in[r]$. Thus by scaling $\alpha_{j}$ by $\mu_{j}$ we may assume that $\sum_{j=1}^{r}\alpha_{j}=0$. Without loss of generality, we may let $H^{+}_{j}=\{x\in\mathbb{R}^{n-d-1}\mid\langle\alpha_{j},x\rangle\geq 0\}$ for all $j\in[r]$.
@@
For each $j\in[r]$, define $\lambda_{j}=(\lambda_{j,1},\ldots,\lambda_{j,n})\in\mathbb{R}^{n}$ by $\lambda_{j,i}=\langle\alpha_{j},g_{i}\rangle$ for all $i\in[n]$. As $\alpha_{j}\neq 0$, by Proposition 2.2 we have that $\lambda_{j}\neq 0$ and that $\sum_{i=1}^{n}\lambda_{j,i}=0$ and $\sum_{i=1}^{n}\lambda_{j,i}a_{i}=0$. Now let $j\in[r]$ and suppose that $i\in I_{j}$. Since $g_{i}\in\operatorname{\mathrm{Int}}(B_{j})$, we have $\lambda_{j,i}>0$ and $\lambda_{k,i}=0$ for all $k\neq j-1,j$. As $\sum_{j=1}^{r}\alpha_{j}=0$, we have $\langle\sum_{j=1}^{r}\alpha_{j},g_{i}\rangle=\langle\alpha_{j-1},g_{i}\rangle+\langle\alpha_{j},g_{i}\rangle=0$. Thus $\lambda_{j-1,i}=-\lambda_{j,i}$ for all $i\in I_{j}$. Again fix $j\in[r]$ and consider the relation $\sum_{i=1}^{n}\lambda_{j,i}a_{i}=0$. If $i\notin I_{j}\sqcup I_{j+1}$, then $g_{i}\notin\operatorname{\mathrm{Int}}(B_{j})\sqcup\operatorname{\mathrm{Int}}(B_{j+1})$. In particular, $g_{i}\in\cap_{\ell\neq k,k-1}H_{\ell}$ for some $k\neq j,j+1$ and so $\langle\alpha_{j},g_{i}\rangle=0$ for all $i\notin I_{j}\sqcup I_{j+1}$. Thus $\sum_{i\in I_{j}}\lambda_{j,i}a_{i}+\sum_{i\in I_{j+1}}\lambda_{j,i}a_{i}=0$ and so $\sum_{i\in I_{j}}\lambda_{j,i}a_{i}=-\sum_{i\in I_{j+1}}\lambda_{j,i}a_{i}=\sum_{i\in I_{j+1}}\lambda_{j+1,i}a_{i}$ for any $j\in[r]$. We therefore have that
(3.2) $\sum_{i\in I_{1}}\lambda_{1,i}a_{i}=\cdots=\sum_{i\in I_{r}}\lambda_{r,i}a_{i}.$
@@ 3.2. Complex Linear Fans
Before giving a complex analogue of Lemma 3.1, we first give a description of complex regular $r$-fans in $\mathbb{C}^{d}$ in terms of the standard Hermitian inner product. Any complex affine hyperplane in $\mathbb{C}^{d}$ is of the form $H_{\mathbb{C}}:=\{z\in\mathbb{C}^{d}\mid\langle\alpha,z\rangle=\beta\}$ for some $(\alpha,\beta)\in\mathbb{C}^{n}\setminus\{0\}\times\mathbb{C}$. Let $\omega_{r}=\exp(2\pi i/r)$ be the standard $r$-th root of unity. If $F_{r}=\cup_{j=1}^{r}B_{j}$ is a complex regular $r$-fan centered about $H_{\mathbb{C}}$, then there exists some unit complex number $\lambda$ such that $B_{j}=\{z\in\mathbb{C}^{d}\mid\langle\alpha,z\rangle=\beta+t\lambda\omega_{r}^{j}\,\,\text{for some}\,\,t\geq 0\}$ for all $j\in[r]$. Moreover, one may take $\lambda=1$ after scaling $(\alpha,\beta)$ by $\lambda^{-1}$.
@@
As before, suppose that $I_{1},\ldots,I_{r}\subset[n]$ are pairwise disjoint non-empty sets such that $\sum_{i\in I_{1}}t_{i}a_{i}=\cdots=\sum_{i\in I_{r}}t_{i}a_{i}$ where $t_{i}>0$ for all $i\in\sqcup_{j=1}^{r}I_{j}$ and $\sum_{i\in I_{j}}t_{i}=1$ for all $j\in[r]$, and moreover that $t_{i}=0$ for all $i\notin\sqcup_{j=1}^{r}I_{j}$. To construct a complex affine dependence among the $a_{i}$, define $\lambda=(\lambda_{1},\ldots,\lambda_{n})\in\mathbb{C}^{n}\setminus\{0\}$ as follows: for each $j\in[r]$ and each $i\in I_{j}$, let $\lambda_{i}=t_{i}\omega_{r}^{j}$, and let $\lambda_{i}=0$ for all $i\notin\sqcup_{j=1}I_{j}$. As $\sum_{i\in I_{j}}t_{j}=1$ for all $j\in[r]$, we have $\sum_{i=1}^{n}\lambda_{i}=\sum_{j\in[r]}\sum_{i\in I_{j}}t_{i}\omega_{r}^{j}=\sum_{j\in[r]}\omega_{j}^{r}=0$. Likewise, we have $\sum_{i\in I_{j}}t_{i}a_{i}=c$ is constant over $j\in[r]$ and so $\sum_{i\in[n]}\lambda_{i}a_{i}=\sum_{j\in[r]}\sum_{i\in I_{j}}t_{i}\omega_{r}^{j}a_{i}=c\sum_{j\in[r]}\omega_{r}^{j}=0$ as well. By Proposition 2.2, there is some $\alpha\neq 0$ such that $\langle\alpha,g_{i}\rangle=\lambda_{i}$ for all $i\in[n]$. Let $H_{\mathbb{C}}=\langle\alpha\rangle^{\perp}$ be the resulting linear complex hyperplane in $\mathbb{C}^{n-d-1}$, and for each $j\in[r]$ let $B_{j}=\{z\in\mathbb{C}^{n-d-1}\mid\langle\alpha,z\rangle=t\omega_{r}^{j}\,\,\text{for some}\,\,t\geq 0\}$. Thus $F_{r}=\cup_{j=1}^{r}B_{j}$ is a complex regular $r$-fan $F_{r}$ centered about $H_{\mathbb{C}}$. By construction, $g_{i}\in\operatorname{\mathrm{Int}}{B_{j}}$ for all $i\in I_{j}$ and any $j\in[r]$, while $g_{i}\in H_{\mathbb{C}}$ if $i\notin\sqcup_{j\in[r]}I_{j}$. ∎
@@ 4. Distributions by a Single Fan
In this section we prove our distribution results for a single fan, each of which follows by applying Gale duality to the linear version of an associated topological Tverberg-type theorem.
4.1. Proof of Theorems 1.1 and  1.3
@@
Let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. We let $n\geq(r-1)(d+m+1)$ when $\mathbb{K}=\mathbb{R}$ and $n\geq(r-1)(2d+m+1)+1$ when $\mathbb{K}=\mathbb{C}$. Suppose that $X=X_{1}\sqcup\cdots\sqcup X_{m}$ is an $m$-coloring of a set of $n$ affinely spanning points in $\mathbb{K}^{n-d-1}$. Corresponding to each $X_{i}$, define a family of non-empty subsets of $X$ by $\mathcal{F}_{i}=\{F\subseteq X\mid|F\cap X_{i}|>|X_{i}|/r\}$ and let $\mathcal{F}=\cup_{i=1}^{m}\mathcal{F}_{i}$. As no $\mathcal{F}_{i}$ can contain $r$ pairwise disjoint members by the pigeon-hole principle, $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$. Theorem 1.4 now guarantees an $r$-fan $F_{r}=\cup_{j=1}^{r}B_{j}$ in $\mathbb{K}^{n-d-1}$ (a conical $r$-fan when $\mathbb{K}=\mathbb{R}$, and a complex regular fan when $\mathbb{K}=\mathbb{C})$ such that no $\operatorname{\mathrm{Int}}(B_{j})$ contains any element of $\mathcal{F}$. It follows that $|\operatorname{\mathrm{Int}}(B_{j})\cap X_{i}|\leq|X_{i}|/r$ for each $j\in[r]$ and $i\in[m]$, so that $F_{r}$ is the desired $r$-fan. Indeed, were $|\operatorname{\mathrm{Int}}(B_{j})\cap X_{i}|>|X_{i}|/r$ for some $j\in[r]$ and $i\in[m]$ then $\operatorname{\mathrm{Int}}(B_{j})\cap X$ would be an element of $\mathcal{F}_{i}$ and so $\operatorname{\mathrm{Int}}(B_{j})\cap X$ would not be a subset of $\operatorname{\mathrm{Int}}(B_{j})$. ∎
4.2. Proof of Theorem 1.4
We now prove our piercing distribution result for a single fan.
@@
Again let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$ and let $X=\{x_{1},\ldots,x_{n}\}$ be a set of $n$ affinely spanning points in $\mathbb{K}^{n-d-1}$, where $n\geq(r-1)(d+m+1)+1$ when $\mathbb{K}=\mathbb{R}$ and $n\geq(r-1)(2d+m+1)+1$ if $\mathbb{K}=\mathbb{C}$. We are given that $\mathcal{F}$ is a family of non-empty subsets of $X$ satisfying $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$. We now lift $X$ to the $\mathbb{K}$-affine hyperplane $\mathbb{K}^{n-d-1}\times\{1\}$ in $\mathbb{K}^{n-d}$. The sequence $g_{1}=(x_{1},1),\ldots,g_{n}=(x_{n},1)$ in $\mathbb{K}^{n-d}$ is linearly spanning, and including $g_{n+1}=-\sum_{i=1}^{n}g_{i}$ gives a sequence of $n+1$ points in $\mathbb{K}^{n-d}$ which both linearly span and which sum to zero. By Proposition 2.1, the sequence $g_{1},\ldots,g_{n+1}$ is the Gale transform of $n+1$ points $a_{1},\ldots,a_{n+1}$ in $\mathbb{K}^{d}$. We let $f\colon\Delta_{n}\rightarrow\mathbb{K}^{d}$ be the (real) linear map determined by sending each $i\in[n+1]$ to $a_{i}$ (here we identify $\mathbb{C}^{d}$ with $\mathbb{R}^{2d}$).
@@
Corresponding to $\mathcal{F}$, let $\mathcal{F}^{\prime}$ be the family of all subsets $\{i_{1},\ldots,i_{k}\}$ of $[n]$ such that $\{x_{i_{1}},\ldots,x_{i_{k}}\}$ is in $\mathcal{F}$. As $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$, $\chi(\mathrm{KG}^{r}(\mathcal{F}^{\prime}))\leq m$ as well. Given the value of $n$, we may apply Theorem 1.5 to the restriction $f\mid_{\Delta_{n-1}}\colon\Delta_{n-1}\rightarrow\mathbb{K}^{d}$ of $f$ to the simplex $\Delta_{n-1}$ whose vertex set is $[n]$. Thus there is a proper Tverberg tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f|_{\Delta_{n-1}}$ such that each $I_{j}=\operatorname{\mathrm{Vert}}(\sigma_{j})\subset[n]$ does not contain any element of $\mathcal{F}^{\prime}$. As $(\sigma_{1},\ldots,\sigma_{r})$ is also a proper $r$-Tverberg partition for the original map $f$, by Lemmas 3.1 and  3.2) there is a linear $r$-fan $F_{r}=\cup_{j=1}^{r}B_{j}$ in $\mathbb{K}^{n-d}$ (conical when $\mathbb{K}=\mathbb{R}$, and complex linear when $\mathbb{K}=\mathbb{C}$) which distributes $g_{1},\ldots,g_{n+1}$ according to the $I_{j}$. Thus $g_{i}\in\operatorname{\mathrm{Int}}(B_{j})$ for all $i\in I_{j}$ and all $j\in[r]$, while $g_{i}\in C=\cap_{j=1}^{r}H_{j}$ for all $i\notin\sqcup_{j=1}^{r}I_{j}$. As $n+1$ does not lie in any of the $I_{j}$, we have that $g_{n+1}$ lies in $C$.